OpenStudy (anonymous):

If you have a 2.50 mole sample of a substance, how many representative particles does this represent? 2.50 x 1023 6.02 x 1023 1.51 x 1024 2.50 x 1024

OpenStudy (xishem):

Once again, we need to use a conversion factor:\[2.50mol*(\frac{6.022 \times {10^{23}}particles}{1mol})=1.51 \times {10^{24}}\ particles\]

OpenStudy (unklerhaukus):

Avogadros numbers is the number of molecule in 1 mole